OpenStudy (anonymous):

2y+3/y = 3/2 rational expression help

OpenStudy (anonymous):

@Ashleyisakitty @poopsiedoodle @whpalmer4 @blahhedup101

OpenStudy (freckles):

are you trying to solve an equation?

OpenStudy (freckles):

\[2y+\frac{3}{y}=\frac{3}{2} \\ \text{ multiply both sides by } y \neq 0 \\ y(2y+\frac{3}{y})=y(\frac{3}{2} )\\ y(2y)+y(\frac{3}{y})=\frac{3}{2}y \\ 2y^2+3=\frac{3}{2} y \\ \text{ if you don't want \to deal with any fractions multiply both sides by 2 } \\ 4y^2+6=3y \\ \text{ subtract} 3y \text{ on both sides } 4y^2-3y+6=0\] you have a quadratic equation to solve. but the equation you wrote is probably not the equation you meant you probably meant \[\frac{2y+3}{y}=\frac{3}{2}\] though I could be wrong about that too you did say rational expression after all and you have an equation composed of rational expressions on left and right side

OpenStudy (freckles):

and if you meant the last equation I wrote which is (2y+3)/y=3/2 I would still multiply both sides by y and 2

OpenStudy (freckles):

\[2(2y+3)=3(y) \\ 4y+6=3y \]

OpenStudy (anonymous):

@freckles than what

OpenStudy (freckles):

you solve for y

OpenStudy (freckles):

And I still don't know what equation you mean

OpenStudy (anonymous):

@freckles The second equation you showed

OpenStudy (freckles):

ok do something to get your y terms on one side

OpenStudy (anonymous):

@freckles Thanks

OpenStudy (freckles):

what did you get?

OpenStudy (freckles):

\[2(2y+3)=3(y) \\ 4y+6=3y\] You subtracted 4y on both sides right?